# Schur’s Positive-Definite Network: Deep Learning in the SPD cone with structure

### v1

@@ 4.1 A model-based formulation: the Graphical Lasso
where $\|\bm{\Theta}\|_{1,\mathrm{off}}$ denote the off-diagonal $\ell_{1}$ norm of $\bm{\Theta}$, equal to $\sum_{i\neq j}|\Theta_{ij}|$. Equation 5 corresponds to an $\ell_{1}$-penalized maximum likelihood estimator under the assumption that the signals are Gaussian with covariance $\bm{\Sigma}$. It can be computed with various convex optimization solvers, e.g. primal formulations (Mazumder and Hastie, 2012), proximal gradient descent (Rolfs et al., 2012), or second order algorithms (Hsieh et al., 2014), all with strict guarantees of recovering both sparse and SPD precision matrices. In particular, the algorithm of Mazumder and Hastie (2012) precisely fits the form of the updates in (2), where the functions $f$ and $g$ are not learnt, but correspond to solutions of changing Lasso regression problems.
Learning-based approaches, however, are scarcer. To the best of our knowledge, DeepGraph (Belilovsky et al., 2017) and GLAD (Shrivastava et al., 2019) are the only two existing works to tackle sparse precision matrix estimation under a data-driven perspective, both with restrictive limitations. This somewhat surprising fact is justified by the challenges posed by the task, which we highlight in the sequel.
4.2 Limitations of learning-based approaches for sparse precision matrix estimation
DeepGraph.Belilovsky et al. (2017) propose to directly map empirical covariance matrices to the graph structures using a CNN-based architecture trained with supervision. DeepGraph is trained for structure recovery only and not for the estimation of the actual precision matrix. It disregards the magnitude of the elements of the output matrix and focuses solely on its support, thus circumventing the problem of SPD learning, which is a limitation for some applications.
@@ 4.2 Limitations of learning-based approaches for sparse precision matrix estimation
Figure 2: GLAD limitations: Smallest eigenvalue (orange), density degree (green) and relative discrepancy of the two matrices (blue) for one of the output $(\bm{Z}_{\mathrm{out}},\bm{\Theta}_{\mathrm{out}})$ of GLAD. $\bm{\Theta}_{\mathrm{out}}$ and $\bm{Z}_{\mathrm{out}}$ are significantly different, $\bm{\Theta}_{\mathrm{out}}$ is not sparse, and $\bm{Z}_{\mathrm{out}}$ is not positive-definite.
GLAD. Shrivastava et al. (2019) unroll an alternating minimization algorithm for the Graphical Lasso (Dalal and Rajaratnam, 2017) to serve as an inductive bias for their model, which learns to predict elementwise regularization parameters of the updates. GLAD is trained with supervision, by minimizing a reconstruction error between the outputs and the ground truth matrix. The unrolled algorithm is designed to solve a Lagrangian relaxation of (5)
As detailed in Section A.1, GLAD jointly operates on two matrices $\bm{Z}$ and $\bm{\Theta}$. GLAD is a powerful model, and has a minimalist yet effective parametrization. Unfortunately, by treating $\bm{Z}$ and $\bm{\Theta}$ separately, the architecture cannot learn matrices that are both sparse and SPD. This is illustrated in Figure 2 where $\bm{\Theta}$ and $\bm{Z}$ differ, $\bm{\Theta}$ is not sparse and $\bm{Z}$ is not SPD.
4.3 New models: SpodNet for learning sparse precision matrices
@@
In this section, we present how to leverage SpodNet layers to overcome the limitations of existing solutions for learning both sparse and SPD precision matrices. By order of complexity, we present three neural architectures relying on SpodNet layers, each corresponding to a specific choice of the functions $f$ and $g$ in Equation 2.
@@
All three architectures are inspired by an unrolling of a proximal block coordinate descent applied to the Graphical Lasso problem (5), where the blocks are column-row pairs described in Section 3. Because (full) proximal gradient descent on (5) is called Graphical ISTA (Rolfs et al., 2012), we coin this Block-Graphical ISTA.
To derive its iterations, we note that the gradient of $\bm{\Theta}\mapsto-\log\det(\bm{\Theta})+\langle\bm{S},\bm{\Theta}\rangle$ (the smooth term in the GLasso loss) is $-\bm{\Theta}^{-1}+\bm{S}$ (Boyd and Vandenberghe, 2004, Section A.4.1) and thus the gradient restricted to the $\bm{\theta}_{12}$-block is $-[\bm{\Theta}^{-1}]_{12}+\bm{s}_{12}=-\bm{w}_{12}+\bm{s}_{12}$. Consequently, the proximal gradient descent algorithm on these coordinates writes $\bm{\theta}_{12}^{+}=\mathrm{prox}_{\lambda\gamma\|\cdot\|_{1}}(\bm{\theta}_{12}-\gamma(\bm{s}_{12}-\bm{w}_{12}))$ where $\gamma>0$ is some stepsize and $\mathrm{prox}$ the proximal operator.
Using that the soft-thresholding is the proximal operator of the $\ell_{1}$ norm, Block-Graphical ISTA hence reads
$\bm{\theta}_{12}^{+}=\operatorname{ST}_{\gamma\lambda}(\bm{\theta}_{12}-\gamma(\bm{s}_{12}-\bm{w}_{12}))\,.$ (7)
@@
First, we propose to learn the stepsizes and soft-thresholding levels when unrolling the Block Graphical ISTA iterations (7). In a learning-based approach, we use as updating function $f$,
$f_{\mathrm{UBG}}:\bm{\theta}_{12}\mapsto\operatorname{ST}_{\Lambda^{+}}(\bm{\theta}_{12}-\gamma^{+}(\bm{s}_{12}-\bm{w}_{12}))\,,$ (8)
where the elementwise soft-thresholding parameters $\Lambda^{+}$ along with the step-size $\gamma^{+}$ are predicted by perceptrons at each update. We emphasize that this is different from hyperparameter tuning, since $\gamma^{+}$ and $\Lambda^{+}$ are predicted at each update instead of being global parameters. UBG exhibits the highest inductive bias among the models, as it’s architecture is directly derived from unrolling an iterative optimization algorithm designed to minimize a model-based loss. Consequently, this architecture is particularly suited for learning sparse precision matrices.
However, as highlighted in Section 3, the true power of SpodNet lies in the fact that one can plug any update functions and still get SPD outputs. This first motivates extending UBG to a more expressive setting with less inductive bias.
@@
Next, we extend UBG to the Plug-and-Play setting (Venkatakrishnan et al., 2013; Romano et al., 2017; Kamilov et al., 2023). In a nutshell, these type of methods replace the proximal operator in first-order algorithms by a denoiser, usually implemented by a neural network. In our context, this corresponds to replacing the soft-thresholding of UBG by an arbitrary network $\Psi:\mathbb{R}^{p-1}\to\mathbb{R}^{p-1}$, that is,
@@
To promote sparsity of the output of $f_{\mathrm{PnP}}$, the last layer of $\Psi$ performs an elementwise soft-thresholding. The parameter for this soft-thresholding is also learned from the data and given by the same perceptron that predicts $\Lambda^{+}$ in UBG.
End-to-end Learnable Block Graphical-ISTA (E2E)
@@
$f_{\mathrm{E2E}}:\bm{\theta}_{12}^{+}\mapsto\Phi(\bm{\theta}_{12})\,,$ (10)
where $\Phi$ takes in the current state of the column $\bm{\theta}_{12}$ and learns to predict an adequate column update $\bm{\theta}_{12}^{+}$. Intuitively, the neural network $\Phi$ acts as learning both the forward and the backward steps of a forward-backward iteration (Combettes and Pesquet, 2011). E2E is the most expressive of our three models because it is free from any inductive bias. As for PNP, sparsity of the predictions is enforced by a soft-thresholding non-linearity in the last layer of $\Phi$.
Figure 3: Potential instabilities with no normalization: Along column update indices, we plot the smallest eigenvalue (blue), largest diagonal value (orange) and conditioning (green) of $\bm{\Theta}$. We observe that $\bm{\Theta}$ remains SPD but gets increasingly closer to being singular during the updates.
Stability and update of the diagonal coefficient $\theta_{22}$
Although the positive definiteness of $\bm{\Theta}$ is guaranteed to be preserved at each update for any positive-valued function $g$, we have empirically observed that its smallest eigenvalue could approach 00 as visible in Figure 3. In practice this leads to instability in training. Below, we offer an interpretation of this phenomenon and suggest a solution to address it. Each column-row update inside a SpodNet layer can be written as the rank-$2$ update
@@
where $\lambda_{k}$ denotes the $k$-th largest eigenvalue of a SPD matrix. Hence, one way to ensure that the perturbation of the spectrum is small is to control $\|\Delta_{\bm{\Theta}}\|_{\mathrm{op}}$, for instance, by limiting the magnitude of the updates. Experimentally, the most successful approach to control $\|\Delta_{\bm{\Theta}}\|_{\mathrm{op}}$ is to limit the magnitude of $\theta^{+}_{22}$, which is not a trivial task as the updated value $\theta^{+}_{22}=g(\bm{\Theta})+\bm{\theta}^{+}_{21}[\bm{\Theta}_{11}]^{-1}\bm{\theta}^{+}_{12}$ must satisfy $\theta^{+}_{22}-\bm{\theta}^{+}_{21}[\bm{\Theta}_{11}]^{-1}\bm{\theta}^{+}_{12}>0$. Thus, in practice we propose to scale $\bm{\theta}_{12}^{+}$, ultimately choosing for all 3 models $f\in\{f_{\mathrm{UBG}},f_{\mathrm{PnP}},f_{\mathrm{E2E}}\}$ the column update
$\bm{\theta}_{12}^{+}=\sqrt{\zeta}\frac{f(\bm{\Theta})}{\sqrt{f(\bm{\Theta})^{\top}[\bm{\Theta}_{11}]^{-1}f(\bm{\Theta})}}\,,$ (13)
which ensures that the diagonal update is given by $\theta_{22}^{+}=g(\bm{\Theta})+\zeta$. The hyperparameter $\zeta>0$ acts as a form of regularization that handles a compromise between stability and expressivity of the model. In all of our experiments we set $\zeta=1$.
Finally, for all models, we use for $g$ a small neural network that takes as input the previous value of $\theta_{22}$ and the corresponding diagonal entry of $\bm{S}$, $s_{22}$, and whose positivity is ensured by using an exponential function as final nonlinearity.
@@ 5 Experiments
We now illustrate our three graph learning models built on SpodNet layers’s ability to learn relevant sparse precision matrices. This section aims to achieve two objectives. First, we highlight that our three models learn matrices that are indeed sparse and SPD in contrast to other learning-based approaches such as GLAD. Second, we show that our models identify accurate sparse precision matrices and even outperform an oracle Graphical Lasso estimator.
Figure 4: Training dynamics of our UBG, PNP and E2E models, compared to the baseline Oracle GLasso (in red) which provides both sparse and SPD matrices.
Data
@@
We use synthetically generated data to fit our models. We generate $N$ sparse SPD matrices of dimension $p=50$ using sklearn’s make_sparse_spd_matrix function (Pedregosa et al., 2011), of which we ensure proper conditioning by adding $0.1$ to the diagonal entries. These matrices are treated as ground truth precision matrices $\bm{\Theta}_{\mathrm{true}}^{(i)}$ for $i\in[N]$. The sparsity degree of each matrix is controlled through the one imposed on their Cholesky factors during their generation. We consider two setups for these sparse SPD matrices: weakly sparse ones ($\alpha=0.7$), and strongly sparse ones ($\alpha=0.95$). From each of these $N$ ground truth precision matrices $\bm{\Theta}_{\mathrm{true}}=\bm{\Sigma}_{\mathrm{true}}^{-1}$, we sample $n=50$ i.i.d. centered Gaussian random vectors $\bm{x}_{j}\sim\mathcal{N}(0,\bm{\Sigma}^{(i)}_{\mathrm{true}}),j\in[n]$, which are used to compute an empirical covariance matrix $\bm{S}^{(i)}$. Both the training and testing datasets thus comprise distinct couples $(\bm{S}^{(i)},\bm{\Theta}^{(i)}_{\mathrm{true}})$, meaning that each $\bm{S}^{(i)}$ stems from a different distribution. Since $n=p$, the $\bm{S}^{(i)}$’s are poorly conditioned and the estimation of $\bm{\Theta}_{\mathrm{true}}^{(i)}$ is typically considered to be a very difficult problem to solve. We generate $N_{\mathrm{test}}=100$ $(\bm{\Theta},\bm{S})$ couples for the testing set on which we evaluate all approaches. The average sparsity degrees of the ground truth in the weakly sparse and strongly sparse testing sets are respectively about 24% and 90%. The training set for our UBG model comprises 100 couples, whereas PNP and E2E are trained on 1000 couples.
Minimization objective & performance metric

### v3

@@ 4.1 Inferring sparse precision matrices
where $\|\bm{\Theta}\|_{1,\mathrm{off}}$ denote the off-diagonal $\ell_{1}$ norm of $\bm{\Theta}$, equal to $\sum_{i\neq j}|\Theta_{ij}|$. It can be efficiently computed (Rolfs et al., 2012; Mazumder and Hastie, 2012; Hsieh et al., 2014) but suffers from the known limitations from the model-based approaches (Adler and Öktem, 2018; Arridge et al., 2019). To circumvent these limitations, two learning-based approaches have been proposed. Belilovsky et al. (2017) introduced DeepGraph, a CNN that directly maps empirical covariance matrices to the graph structures (i.e. to the support of the precision matrix). We emphasize that this model does not estimate $\bm{\Theta}$ but only its support. Shrivastava et al. (2020) proposed GLAD, a neural architecture based on unrolling an alternating minimization algorithm for solving the GLasso. Based on a Lagrangian relaxation, GLAD iterates over two matrices $\bm{Z}$ and $\bm{\Theta}$ while learning step-size and thresholding parameters (see Section B.2 for more details). Out of these two matrices $\bm{Z}$ is sparse and $\bm{\Theta}$ is SPD, but GLAD fails to learn a single matrix that is both sparse and SPD. This limitation is highlighted in Figure 2 showing that, during training, $\bm{\Theta}$ and $\bm{Z}$ differ, $\bm{\Theta}$ is not sparse and $\bm{Z}$ is not SPD.
@@
Figure 2: GLAD’s limitations. Smallest eigenvalue (orange), density degree (green) and relative discrepancy of the two matrices (blue) for an output $(\bm{Z}_{\mathrm{out}},\bm{\Theta}_{\mathrm{out}})$ of GLAD. $\bm{\Theta}_{\mathrm{out}}$ and $\bm{Z}_{\mathrm{out}}$ are different, $\bm{\Theta}_{\mathrm{out}}$ is not sparse, and $\bm{Z}_{\mathrm{out}}$ is not positive-definite.
4.2 SpodNet for sparse SPD learning
@@
We now show how SpodNet overcomes the limitations of current methods for learning both sparse and SPD matrices. By order of complexity, we introduce three new neural architectures, each corresponding to a specific choice of the functions $f,g$ for learning the values of $\bm{u}$ and $v$ in Equation 2.
@@
Our choices for $f,g$ are inspired by an unrolling of a proximal block coordinate descent applied to the GLasso problem, where the blocks are column-row pairs described in Section 3. A proximal coordinate gradient descent step on the GLasso objective updates the value of $\bm{\theta}_{12}$ with
$\bm{\theta}_{12}^{+}=\operatorname{ST}_{\gamma\lambda}(\bm{\theta}_{12}-\gamma(\bm{s}_{12}-\bm{w}_{12}))\,,$ (6)
where $\gamma>0$ is a step-size. This stems from the fact that the gradient of $\bm{\Theta}\mapsto-\log\det(\bm{\Theta})+\langle\bm{S},\bm{\Theta}\rangle$is $-\bm{\Theta}^{-1}+\bm{S}$ (Boyd and Vandenberghe, 2004, Section A.4.1), which restricted to the $\bm{\theta}_{12}$-block gives $-[\bm{\Theta}^{-1}]_{12}+\bm{s}_{12}=-\bm{w}_{12}+\bm{s}_{12}$. Together with the fact that the proximal operator of the $\ell_{1}$ norm is the soft-thresholding (Parikh et al., 2014) we get Equation 6. Because (full) proximal gradient descent on the GLasso is called Graphical ISTA (Rolfs et al., 2012), we coin Equation 6 Block-Graphical ISTA. We can now introduce our three architectures.
@@
First, we propose to learn the stepsizes and soft-thresholding levels when unrolling the Block Graphical ISTA iterations Equation 6. In a learning-based approach, we use as updating function $f$,
$f_{\mathrm{UBG}}:\bm{\theta}_{12}\mapsto\operatorname{ST}_{\bm{\lambda}^{+}}(\bm{\theta}_{12}-\gamma^{+}(\bm{s}_{12}-\bm{w}_{12}))\,,$ (7)
where the step-size $\gamma^{+}=\operatorname{NN}_{1}(\bm{\theta}_{12})>0$ and the soft-thresholding parameters $\bm{\lambda}^{+}=\operatorname{NN}_{2}(\bm{\theta}_{12}-\gamma^{+}(\bm{s}_{12}-\bm{w}_{12}))\in\mathbb{R}^{p-1}$. $\operatorname{NN}_{1}$ and $\operatorname{NN}_{2}$ are small multilayer perceptrons with architectures provided in Section A.2. We emphasize that this is different from hyperparameter tuning, since $\gamma^{+}$ and $\bm{\lambda}^{+}$ are predicted at each update instead of being global parameters. UBG exhibits the highest inductive bias among the models, as its architecture is directly derived from unrolling an iterative optimization algorithm designed to minimize a model-based loss.
As highlighted in Section 3, one can plug any update functions and still get SPD outputs. This motivates extending UBG to a more expressive architecture.
@@
Precisely, we extend UBG to a Plug-and-Play-like setting (Venkatakrishnan et al., 2013; Romano et al., 2017). In a nutshell, these methods replace the proximal operator in first-order algorithms by a denoiser, usually implemented by a neural network. In our context, this corresponds to replacing the soft-thresholding of UBG by a neural network $\Psi:\mathbb{R}^{p-1}\to\mathbb{R}^{p-1}$, that is,
@@
To promote sparsity of the output of $f_{\mathrm{PnP}}$, the last layer of $\Psi$ performs an elementwise soft-thresholding. The parameter for this soft-thresholding is also learned from data and given by the same multilayer perceptron that predicts $\bm{\lambda}^{+}$ in UBG (Section A.2).
End-to-end updates (E2E)
@@
$f_{\mathrm{E2E}}:\bm{\theta}_{12}\mapsto\Phi(\bm{\theta}_{12})\,,$ (9)
where $\Phi$ takes in the current state of the column $\bm{\theta}_{12}$ and learns to predict an adequate column update $\bm{\theta}_{12}^{+}$. Intuitively, the neural network $\Phi$ acts as learning both the forward and the backward steps of a forward-backward iteration (Combettes and Pesquet, 2011). As for PNP, sparsity of the predictions is enforced by a soft-thresholding non-linearity in the last layer of $\Phi$.
Figure 3: Training dynamics of our 3 models (on test data described in Section 5.1). Left: The outputs of our 3 models remain positive-definite. Middle: The conditioning remains stable. Right: The outputs are sparse. Overall our models produce jointly sparse + SPD outputs.
Finally, for all models, we use for $g$ a small neural network that takes as input $\theta_{22},s_{22}$ and $(\bm{\theta}_{12}^{+})^{\top}[\bm{\Theta}_{11}]^{-1}\bm{\theta}_{12}^{+}$. Its positivity is ensured by using an absolute value function as final nonlinearity.
Improving training stability
Although the positive definiteness of $\bm{\Theta}$ is guaranteed to be preserved at each update for any positive-valued function $g$, we have empirically observed that its smallest eigenvalue could approach 00 as visible in Section B.1. In practice this leads to instability in training. Below, we provide an interpretation of this phenomenon and provide a solution to address it. Each column-row update inside a SpodNet layer can be written as the rank-$2$ update
@@
where $\lambda_{k}$ denotes the $k$-th largest eigenvalue of a SPD matrix. Hence, one way to ensure that the perturbation of the spectrum is small is to control $\|\Delta_{\bm{\Theta}}\|_{\mathrm{op}}$, for instance, by limiting the magnitude of the updates. Experimentally, the most successful approach to control $\|\Delta_{\bm{\Theta}}\|_{\mathrm{op}}$ is to limit the magnitude of $\theta^{+}_{22}$, which is not a trivial task as the updated value $\theta^{+}_{22}=g(\bm{\Theta})+\bm{\theta}^{+}_{21}[\bm{\Theta}_{11}]^{-1}\bm{\theta}^{+}_{12}$ must satisfy $\theta^{+}_{22}-\bm{\theta}^{+}_{21}[\bm{\Theta}_{11}]^{-1}\bm{\theta}^{+}_{12}>0$. Thus, we propose to scale $\bm{\theta}_{12}^{+}$ by scaling the preactivation [^1] Experimentally scaling the preactivation works better than scaling the network’s output. $\bm{z}$ of the last layer of $f$ by $\frac{\sqrt{\zeta}}{\sqrt{\bm{z}^{\top}[\bm{\Theta}_{11}]^{-1}\bm{z}}}$. The hyperparameter $\zeta>0$ acts as a form of regularization that handles a compromise between stability and expressivity of the model; in all of our experiments we use $\zeta=1$. As visible in Figure 3 this scaling ensures a smooth training of the network, with stable condition number of the $\bm{\Theta}$ matrix.
@@ 5 Experiments
We now illustrate SpodNet’s ability to learn SPD matrices with additional structure by using our three derived graph learning models (UBG, PNP and E2E) to learn sparse precision matrices, on both synthetic and real data. Additional details regarding the experimental setups can be found in Appendix A. PyTorch implementations are provided with the submission and will be made public upon acceptance.
5.1 Sparse precision matrix recovery
Our goal in this section is to evaluate our models’ performance and generalization ability regarding the recovery of sparse precision matrices on synthetic data. The following experiments serve several purposes: (1) compare learning-based methods against traditional model-based methods, (2) evaluate our models’ reconstruction performance in terms of matrix estimation and support recovery.
Data & training
@@ 5.1 Sparse precision matrix recovery
We use synthetic data to train our SpodNet models as in Belilovsky et al. (2017); Shrivastava et al. (2020). We generate $N$ sparse SPD $p\times p$ matrices using sklearn’s make_sparse_spd_matrix function (Pedregosa et al., 2011), of which we ensure proper conditioning by adding $0.1\cdot\mathbf{I}_{p}$. These matrices are treated as ground truth precision matrices $\bm{\Theta}_{\mathrm{true}}^{(i)}$ for $i\in[N]$. The sparsity degree of each matrix is controlled through the one imposed on their Cholesky factors during their generation: in the strongly sparse setting $\bm{\Theta}_{\mathrm{true}}$ has roughly 90 % of zero entries while in the weakly sparse setting this value is around 25 %. For each of these $N$ ground truth precision matrices $\bm{\Theta}_{\mathrm{true}}$, we sample $n$ i.i.d. centered Gaussian random vectors $\bm{x}_{j}\sim\mathcal{N}(0,(\bm{\Theta}^{(i)}_{\mathrm{true}})^{-1})$, which are used to compute an empirical covariance matrix $\bm{S}^{(i)}$. The dataset thus comprises couples $(\bm{S}^{(i)},\bm{\Theta}^{(i)}_{\mathrm{true}})$ where each $\bm{S}^{(i)}$ stems from a different ground truth precision matrix. For all the experiments, we generate $N_{\mathrm{train}}=1000$ different $(\bm{S},\bm{\Theta}_{\mathrm{true}})$ couples for the training set and $N_{\mathrm{test}}=100$ couples for the testing set on which we validate our models. Further details on the parameters used during the data generation are in Section A.1.